Hi, I have been having trouble on my math homework on this question and I was wondering if you could please help me: If the exterior angles of a convex octagon are x+13, 2x-1, 2x+6, 2x+12, 2x-17, 3x-4, 3x-10, and 4x, calculate the smallest of the eight angles.
I would really appreciate your help, I've been stuck on this problem all day. Thanks for taking your time.

Answer by stanbon(48558)

If the exterior angles of a convex octagon are x+13, 2x-1, 2x+6, 2x+12, 2x-17, 3x-4, 3x-10, and 4x, calculate the smallest of the eight angles.
-------------------
The exterior angles of a convex polygon always add up to 360 degrees.
EQUATION:
Adding your angles you get: 19x-14 = 360
19x = 374
x = (374/19) degrees
-------------------
2x-1 is the smallest angle:
2(374/19) -1= 38.3684 degrees
================================
Cheers,
Stan H.